for the variable triangle ABC with fixed vertex C(1,2) and A,B having coordinates (cost,sint) , (sint,-cost) respectively , find the locus of its centroid

Grade:

1 Answers

jitender lakhanpal
12 years ago

Dear krishnan,

coordinates of orthocentre for this triangle (cost +sint +1)/3 , (sint - cost +2 )/3

h = (cost +sint +1)/3     k =   (sint - cost +2 )/3

 

(3h-1)2= (cost +sint)2 ---1     (3k-2)2 = (sint - cost)2---2             

now squaring 1 and 2 and adding them we get

9(h2+k2)-6h -12k +3 = 0 

now replacing h , k by x, y

9(x2+y2)-6x -12y +3 = 0   is the locus of the coordinates of centeroid 

 

 

 

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar: Click here to download the toolbar..  

jitender 

askiitians expert